A concrete pipe of diameter 1 000 mm is laid at a slope of 1 in 100. Estimate the rate at which the water will run down the pipe so that it remains uniformly half full. Use C = 49 SI Units.


1
Expert's answer
2021-11-17T07:14:01-0500

Half Full Circular concrete pipe-

Diameter (D) = 1000 mm = 1 m

Slope (S) = 1 in 100

Chezy's Constant (C) = 49

To Find-

The rate at which the water will run down the pipe (Q).

"Discharge(Q)=AC(RS)^{1\/2}"

"A=\pi/4 (D)^2 =0.3927 m^2"

"C=49"

"R=\text{hydraulic mean depth} =A/P"

"P=perimeter=\pi D/2"

"R=A\/P \u2014\u2014> =D\/4"

"S=1\/100"

Now,

"Q=0.3927\times49\times (1/4 \times 1/100)^{1/2}"

"Q=0.962m^3/s \space (\text{this is the final answer})"

The rate at which the water will run down the pipe (Q) = 0.962 m3/s